It took a man 3 seconds nto raise a load of 40kg to a hight of 25 m.Determine the average power employed.

1
Expert's answer
2022-02-06T14:32:18-0500

Given:

"t=3\:\rm s"

"m=40\:\rm kg"

"h=25\:\rm m"

"g=9.8\:\rm N/kg"


The average power

"P=\frac{W}{t}=\frac{mgh}{t}"

"P=\frac{40*9.8*25}{3}=3267\:\rm W"